# M2OST: Many-to-one Regression for Predicting Spatial Transcriptomics from Digital Pathology Images

### v2

@@ Introduction
Due to the inherently sparse nature of the ST map, we need to filter out less-variable genes in each dataset before initiating the training process. Following previous works (He et al. 2020), we select 250 spatially variable genes for the regression task, which will be released with the code.
Using the same pre-processing procedures in (Zeng et al. 2022), we normalize the gene expression counts for each spot by dividing them by their total counts and multiplying the result by the scale factor 1,000,000. Then, the values are natural-log transformed by log(1 + $x$) where $x$ is the normalized count. Given the high sparsity of ST maps, we select 250 spatially variable genes for regression on each dataset.
@@
Then, we substitute the three ITMMs into one unified Self-Attention to directly process the concatenated sequences (the three sequences are of the same length without DPE, so they can be directly concatenated), destroying the decoupled design in M2OST. It is observed that the parameter count dramatically increased, but the model performance did not benefit from it, which validates the efficiency of using ITMM to decouple the multi-scale feature extraction process in M2OST. We further remove CTMM from M2OST, using simple concatenation for cross-level feature fusion. This time, the parameter count did not drop much, while the performance suffered a further decline. This indicates that CTMM is necessary for processing such many-to-one modeling problems, where each sequence may contain different semantic information that cannot be fused by simple concatenation. We have also tried using summation to replace CTMM, but it fails to outperform the concatenation scheme.
@@
Figure 7: (a) Visualization of the ST map after PCA. (b) Visualization of the spatial distribution of the BRD2 gene.
@@
The experimental results of the comparison between M2OST and other methods are presented in Table 3. This table provides detailed insights into the PCC and RMSE on various datasets of different methods, along with their parameter count and FLOPs. Analysis of the experimental results reveals that M2OST achieves superior performance with fewer FLOPs and a reduced parameter count. In comparison to ST-Net, which features 0.38 million more parameters and 1.43G more FLOPs, M2OST surpasses its performance on HER2+ and cSCC datasets by 0.97% and 1.23% PCC, respectively.
@@ Appendix A Selected Genes
Table 5: Performance of M2SOT after masking one random level of image in different percentages of the samples in the fused dataset (i.e., the dataset combining HBC, HER2+ and cSCC together).
@@ Appendix B Additional Experimental Results
The experimental results are presented in Table 6. As shown, using these backbones for patch embedding leads to further performance improvement compared with the original M2SOT. On one hand, this indicates that these existing backbones are able to generate more accurate token representations for M2OST, and on the other hand, it also demonstrates that many-to-one modeling can further improve the performance of these backbones. In our future works, we will investigate how to more efficiently introduce the many-to-one scheme to existing backbones.

### v3

@@ Introduction
Given the inherently sparse nature of the ST map, we filter out less-variable genes in each dataset based on the criteria outlined in (He et al. 2020), eventually preserving 250 spatially variable genes per dataset for training. As for the pre-processing procedures, they are also kept identical to those described in (He et al. 2020). Specifically, we normalize the gene expression counts for each spot by dividing them by the sum of expressions within that spot, then multiplying the result by a scale factor of 1,000,000. The normalized values are subsequently transformed using the natural logarithm, calculated as log(1 + $x$), where $x$ is the normalized count.
@@
Then, we substitute the three ITMMs into one unified Self-Attention to directly process the concatenated sequences (the three sequences are of the same length without DPE, so they can be directly concatenated), destroying the decoupled design in M2OST. It is observed that the parameter count dramatically increased, but the model performance did not benefit from it, which validates the efficiency of using ITMM to decouple the multi-scale feature extraction process in M2OST. We further remove CTMM from M2OST, using simple concatenation for cross-level feature fusion. This time, the parameter count did not drop much, while the performance suffered a further decline. This indicates that CTMM is necessary for processing such many-to-one modeling problems, where each sequence may contain different semantic information that cannot be fused by simple concatenation. We have also tried using summation to replace CTMM, but it even fails to outperform the concatenation scheme.
@@
Figure 7: (a) Visualization of the ST map after PCA. (b) Visualization of the spatial distribution of the DDX5 gene.
@@
The experimental results of the comparison between M2OST and other methods are presented in Table 3. This table provides detailed insights into the PCC and RMSE on various datasets of different methods, along with their parameter count and FLOPs. Analysis of the experimental results reveals that M2OST achieves superior performance with fewer FLOPs and a reduced parameter count. In comparison to ST-Net, which features 0.40M more parameters and 0.63G more FLOPs, M2OST surpasses its performance on HER2+ and cSCC datasets by 1.16% and 1.13% PCC, respectively.
@@ Conclusion
Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2022YFC2504605). It was also supported in part by the Grant in Aid for Scientific Research from the Japanese Ministry for Education, Science, Culture and Sports (MEXT) under the Grant No. 20KK0234, 21H03470.
@@ Appendix A Selected Genes
Table 5: Performance of M2OST after masking one random level of image in different percentages of the samples in the fused dataset (i.e., the dataset combining HBC, HER2+ and cSCC together).
@@ Appendix B Additional Experimental Results
The experimental results are presented in Table 6. As shown, using these backbones for patch embedding leads to further performance improvement compared with the original M2OST. On one hand, this indicates that these existing backbones are able to generate more accurate token representations for M2OST, and on the other hand, it also demonstrates that many-to-one modeling can further improve the performance of these backbones. In our future works, we will investigate how to more efficiently introduce the many-to-one scheme to existing backbones.